OpenStudy (anonymous):

Simplify the product 2p(-3p^2 + 4p - 5)

OpenStudy (anonymous):

@Jhannybean

OpenStudy (unklerhaukus):

\[2p(-3p^2 + 4p - 5)\\=2p\times(-3p^2)+2p\times(4p)+2p\times(-5)\]

OpenStudy (unklerhaukus):

im not that old.

OpenStudy (anonymous):

im not great at simplifying and whatnot so can you finish the problem>.>

OpenStudy (unklerhaukus):

lets look at the first term \[2p×(−3p^2)=(2\times-3)p^{1+2}\] can you simplify this further @Nickrivertit ?

OpenStudy (anonymous):

id say -6^3 + 8p - 10 but thats just guessing really

OpenStudy (unklerhaukus):

that is close,

OpenStudy (unklerhaukus):

the first term simplifies to \[2p×(−3p^2)=(2×−3)p^{1+2}=-6p^3\]

OpenStudy (unklerhaukus):

tell me what \[2p×(4p)\]simplifies to

OpenStudy (anonymous):

no idea, never done math. 8p?

OpenStudy (jhannybean):

\[2p\times(-3p^2)+2p\times(4p)+2p\times(-5)\]\[\large (2*-3)p^{2+1}+(2*4)p^{1+1}+(-5*2)p\]

OpenStudy (unklerhaukus):

\[p\times p =p^2\] so \[2p\times4p=2\times4\times p\times p=8p^2\]

OpenStudy (anonymous):

so is it -6^3 + 8p^2 - 10 or 10p

OpenStudy (unklerhaukus):

\[2p\times-5=-10p\]